OpenStudy (kamille):

of which value of "a" equation "3x^2+ax+24=0" has a solution of 3 . Any help?

OpenStudy (campbell_st):

substitute x = 3 into the equation \[3(3)^2 + a(3) + 24 = 0\] just solve for a

OpenStudy (kamille):

Ah,thank you very much. It is so easy : ) I thought it should be much harder, like "3x^2+ax+24=3(x-3)(x-x1) Thanks!

OpenStudy (campbell_st):

nope is x = 3 is a solution, then it means that it will make the equation = 0... there are 2 values that make it = 0.... but if you had to do that... then you would need to solve by factorising.

OpenStudy (kamille):

Yeah. thank you again :)